You are thinking of investing $3600 this year. You have received advice from family members . Aunt Anne recommends investing in the stock market with a 9.45% average rate of return. . Uncle Rick recommends investing in a 5.60% certificate of deposit (CD). . Grandpa recommends investing in a 0.53% savings account. . Cousin Lisa recommends investing in a 2, 12% money market fund. How much money will you have at the end of 10 years if you pick Aunt Annes advice? money after 10 years: S How much money will you have at the end of 20 years if you pick Uncle Ricks advice? money after 20 years: $ How much money will you have at the end of 40 years if you pick Grandpas advice? money after 40 years: SHow much money will you have at the end of 30 years if you pick your cousin Lisas advice? money after 30 years: S
Answer #1

We use the rule Future Value = Present Value x (1 + rate)^time

Money as a result of Aunt Anne advice: 3600 x (1 + 9.45%)^10 = $8880.97

Money as a result of Uncle Rick advice: 3600 x (1 + 5.60%)^20 = $10,704.86

Money as a result of Grandpa advice: 3600 x (1 + 0.53%)^40 = $4472.64

Money as a result of Lisa advice: 3600 x (1 + 2.12%)^30 = $6,755